# BECoTTA: Input-dependent Online Blending of Experts for Continual Test-time Adaptation

### v2

@@ Abstract
Continual Test Time Adaptation (CTTA) is required to adapt efficiently to continuous unseen domains while retaining previously learned knowledge. However, despite the progress of CTTA, forgetting-adaptation trade-offs and efficiency are still unexplored. Moreover, current CTTA scenarios assume only the disjoint situation, even though real-world domains are seamlessly changed. To tackle these challenges, this paper proposes BECoTTA, an input-dependent yet efficient framework for CTTA. We propose Mixture-of-Domain Low-rank Experts (MoDE) that contains two core components: (i) Domain-Adaptive Routing, which aids in selectively capturing the domain-adaptive knowledge with multiple domain routers, and (ii) Domain-Expert Synergy Loss to maximize the dependency between each domain and expert. We validate our method outperforms multiple CTTA scenarios including disjoint and gradual domain shits, while only requiring $\sim$98fewer trainable parameters. We also provide analyses of our method, including the construction of experts, the effect of domain-adaptive experts, and visualizations. [^1] Project Page: https://becotta-ctta.github.io/.
@@ 1 Introduction
Figure 1: Our BECoTTA achieves superior performance and parameter/memory efficiency against strong CTTA baselines on the CDS-hard scenario. WAD represents the Weather-Augmented Dataset described in  Sec. 3.1
Figure 1: Our BECoTTA achieves superior performance and parameter/memory efficiency against strong CTTA baselines on the CDS-hard scenario. WAD represents the Weather-Augmented Dataset described in  Sec. 3.1
@@
Figure 2: Comparison of TTA process with other SoTA models. We compare the existing models (Wang et al., 2020, 2022a; Song et al., 2023) and denote activated modules as yellow during CTTA process. In particular, CoTTA adopts the mean-teacher architecture and updates the entire model. TENT (Wang et al., 2020) and EcoTTA (Song et al., 2023) update a few parts of the model, however, they achieve suboptimal performance with forgetting. Meanwhile, our BECoTTA updates only MoDE layers for efficient and rapid adaptation while preserving previous knowledge.
@@
Beyond the isolated transferability of traditional TTA approaches on the stationary target domain, Continual Test-Time Adaptation (CTTA) (Niu et al., 2023; Song et al., 2023; Lee et al., 2023a) has been increasingly investigated in recent years, whose goal is to continuously adapt to multiple unseen domains arriving in sequence. Solving the problem of CTTA is crucial because it is closely related to real-world scenarios. For example, let us assume that a vision model in an autonomous vehicle aims to understand road conditions and objects, including pedestrians, vehicles, traffic signs, etc. The agent will encounter different environments over time, according to the changes in the weather, time of the day, and location. Then, the model should rapidly and continuously adapt to these unseen environments while retaining the domain knowledge learned during adaptation as it may re-encounter prior domains in the future.
@@
Therefore, continual TTA approaches need to address the following key challenges: (i) forgetting-adaptation trade-off: retaining previous domain knowledge while learning new domains often limits the model’s plasticity, hindering its ability to learn and adapt to new data, and (ii) computational efficiency: since CTTA models are often assumed to be embedded in edge devices, efficient adaptation is significant. However, as shown in Fig. 2, existing CTTA methods overlook computational efficiency by updating heavy teacher and student models (Wang et al., 2022a) or achieved suboptimal convergence due to updating only a few parts of modules (Wang et al., 2020; Niu et al., 2023; Gan et al., 2023; Gao et al., 2022b).
@@
To tackle these critical issues, this paper proposes Input-dependent Online Blending of Experts for Continual Test-Time Adaptation (BECoTTA) by introducing a surprisingly efficient yet effective module, named Mixture-of-Domain Low-rank Experts (MoDE), atop each backbone block. Our BECoTTA method consists of two key components: (i) Domain Adaptive Routing and (ii) Domain-Expert Synergy Loss. We first propose Domain Adaptive Routing that aims to cluster lightweight low-rank experts (i.e., MoDE modules) with relevant domain knowledge. Next, based on the assignment of domain adaptive routers, we propose Domain-Expert Synergy Loss to maximize the mutual information between each domain and its corresponding expert. In the end, we facilitate cooperation and specialization among domain experts by ensuring strong dependencies. Our modular design allows for selective updates of multiple domain experts, ensuring the transfer of knowledge for each specific domain while preserving previously acquired knowledge. This approach also significantly improves memory and parameter efficiency through sparse updates.
@@
Furthermore, existing CTTA scenarios assume a disjoint change of test domains, where the model encounters a static domain per time step, but do not consider a gradual shift of domains, which is more common in the real world (e.g., seamless weather change like cloudy $\rightarrow$ rainy or afternoon $\rightarrow$ night). To further consider this realistic scenario, we additionally propose Continual Gradual Shifts (CGS) benchmark for CTTA, where the domain gradually shifts over time based on domain-dependent sampling distribution, as illustrated in Fig. 3 top left. This scenario is more advanced than an existing CTTA problem as it demands the model to appropriately adapt each of the input instances, without relying on any implicit guidance from the dominant domain over a given time interval.
@@
We compare our proposed method with strong baselines, including SAR (Niu et al., 2023), DePT (Gao et al., 2022b), VDP (Gan et al., 2023), and EcoTTA (Song et al., 2023), on multiple CTTA scenarios and our suggested CGS benchmark. Our BECoTTA achieves +2.1%p and +1.7%p IoU enhancement respectively on CDS-Hard and CDS-Easy scenarios, by utilizing 95% and 98% fewer parameters used by CoTTA (Wang et al., 2022a). In addition, for the WAD scenario where the source models were additionally pre-trained on the Weather-Augmented Datasets, our model consistently shows +16.8%p performance gain over EcoTTA, utilizing a similar number of parameters on the CDS-Hard scenario (Please see Fig. 2).
@@ 2 Related Works
Continual Test-Time Adaptation (CTTA) (Wang et al., 2022a; Gan et al., 2023; Niu et al., 2023; Song et al., 2023) assumes that target domains are not fixed but change continuously in an online manner. TENT (Wang et al., 2020) is one of the pioneering works, which activates only BatchNorm layers to update trainable affine transform parameters. CoTTA (Wang et al., 2022a) introduces a teacher-student framework, generating pseudo labels from the teacher model and updating it using consistency loss. EcoTTA (Song et al., 2023) utilizes meta networks and self-distilled regularization while considering memory efficiency. DePT (Gao et al., 2022b) plugs visual prompts for efficiently adapting target domains and bootstraps the source representation. However, existing methods often suffer from subordinate convergence since they rely on a shared architecture for adapting test data without considering the correlation between different domains. On the other hand, our BECoTTA introduces a modularized MoE-based architecture where each expert captures domain-adaptive knowledge, and the model transfers only a few related experts for the adaptation of a new domain.
Moreover, recent works (Song et al., 2023; Niu et al., 2022; Lim et al., 2023; Choi et al., 2022a; Liu et al., 2021; Adachi et al., 2022; Jung et al., 2023; Lee et al., 2023b) allow for a slight warmup using the source dataset before deploying the model to CTTA scenario. In particular, TTA-COPE (Lee et al., 2023b) conducts pretraining with labeled source datasets in a supervised manner. EcoTTA (Song et al., 2023) also proceed warm-up to initialize their meta-network. Note that all of them are operationalized in source-free at test time and there is agreement that this setting adheres to the assumptions of CTTA.
@@
Mixture-of-Experts (MoE) (Shazeer et al., 2017; Fedus et al., 2022; Zuo et al., 2021; Wang et al., 2022b) introduces $N$ parallel experts consisting of the feed-forward network with router modules, and sparsely activate a few experts based on their sampling policies. Adamix (Wang et al., 2022b) introduces efficient fine-tuning with Mixture-of-Adaptations structures to learn multiple views from different experts. THOR (Zuo et al., 2021) proposes a new stochastic routing function to prevent inefficiency with routers. Similarly, Meta DMoE (Zhong et al., 2022) adopts the MoE architecture as a teacher model and distills their knowledge to unlabeled target domains, but they do not consider continuous adaptation. In short, to the best of our knowledge, the feasibility of MoE structures is underestimated in the field of CTTA.
@@
Recently, a few continual learning appraoches (Koh et al., 2021; Bang et al., 2021, 2022; Aljundi et al., 2019) have discussed Blurry Continual Learning (Blurry-CL) to better reflect real-world scenarios, beyond the standard CL setting. Blurry-CL assumes that for each sequential task, a majority class exists, and other classes outside the majority class may also overlap and appear. The most renowned scenario setup is a Blurry-M (Aljundi et al., 2019), where the majority class occupies 100-M%, and the remaining classes are randomly composed of M%. While this benchmark handles an overlapping situation, it may not cover practical situations where the domain evolves gradually in CTTA. Therefore, we propose a new benchmark that simulates real-world continual TTA scenarios with a gradual change of domains over time.
@@ 3 Input-dependent Online Blending of Experts for Continual Test-time Adaptation
We first define the problem statement of Continual Test-Time Adaptation (CTTA) in Sec. 3.1. Next, we introduce a novel proposed CTTA method, BeCoTTA, containing Mixture-of-Domain Low-rank Experts (MoDE) and domain-expert synergy loss in Sec. 3.2. After that, we describe the overall optimization process during CTTA in Sec. 3.3.
@@ 3.1 Problem Statement
Continual Test-time Adaptation (CTTA) aims to adapt the pre-trained source model $\bm{f_{s}}$ to continuously changing target domains, as formulated to a task sequence $\bm{X_{t}}=\{X_{t}^{1},X_{t}^{2},..X_{t}^{c},\cdots\}$. The main assumption of CTTA includes two folds: (i) we should not access the source dataset after deploying model to the test time, and (ii) adaptation needs to be done online and in an unsupervised manner. For semantic segmentation tasks, the objective of CTTA is to predict the softmax output $\hat{y_{c}}=f_{c}(\bm{x}_{t}^{c})$ at target domain $c$. $\bm{x}_{t}^{c}$ is sampled from $X_{t}^{c}$, which will be represented $\bm{x}$ from the following sections for brevity.
Figure 3: The overview of BECoTTA. We propose a novel CTTA framework for dynamic real-world scenarios, including disjoint and gradual shifts of domains. When the model receives a target domain input $\bm{x_{t}}$ at timestep $t$, the Domain Discriminator (DD) first estimates a pseudo domain label $d$. Based on the estimated pseudo labels, the domain router $G_{d}$ processes the input to specific experts containing domain-specific information via minimizing our proposed Domain-Expert Synergy Loss $\Theta(D;A)$. Finally, we obtain domain-adaptive representation $h_{d}(\bm{x})$, addressing downstream tasks in test-time.
Parameter Initialization for CTTA. Most CTTA methods use a pre-trained frozen backbone, which contains domain bias from the source domain. This bias impedes the effective transfer of domain-adaptive knowledge in continuous scenarios, due to the predominance of the source domain. To mitigate this limitation, we define the various candidate $D$ domains (e.g., brightness, darkness, blur) and apply the augmentation to the source domain similar to EcoTTA (Song et al., 2023) and DDA (Gao et al., 2022a), coined Weather-Augmented Dataset (WAD). For the augmentation, we utilize pre-trained style-transfer (Jiang et al., 2020) or simple transformations (Buslaev et al., 2020). Through WAD, we acquire domain-specific knowledge before deploying TTA. Note that this process is only done once when setting the source domain. More details are in the Appendix (Sec. D).
@@ 3.2 Mixture-of-Domain Low-rank Experts (MoDE)
The design of Low-rank Experts. For the efficient process during CTTA, we adopt the Sparse Mixture-of-Experts (SMoE) module with a top-k routing policy (Shazeer et al., 2017). Each MoE layer consists of the router $G$ and a set of $N$ lightweight experts, $A_{1}$, $A_{2}$, …, $A_{N}$, where each $A_{i}$ is parameterized by $\bm{W}_{i}^{down}\in\mathbb{R}^{dim\times r}$ and $\bm{W}_{i}^{up}\in\mathbb{R}^{r\times dim}$. Here, $r$ denotes the rank, and $dim$ denotes the embedding dimension of each ViT block. If the $A_{i}$ is activated, it maps the input $\bm{x}$ into the low dimensional space through the projection operation with $\bm{W}_{i}^{down}$. Next, after regularizing the features with a non-linear activation function $\sigma(\cdot)$, it recovers the features to the original dimension using $\bm{W}_{i}^{up}$:
@@
Based on the $G_{d}(\bm{x})$, we selectively update the activated experts associated with the specific domain, inherently isolating them from irrelevant domain knowledge while adapting new ones. In the end, the output of MoDE layer $h_{d}(x)$ aggregate the domain-adaptive features as follows:
@@
This trainable clustering approach allows the model to activate its own set of experts, who are specialized in specific domain knowledge. Moreover, our multi-router-based design accelerates adaptation to the current domain, avoiding interference from knowledge transfer of unrelated domain features. Finally, we perform the skip connection operation with original input $\bm{x}$: $\bm{x}\leftarrow\bm{x}+h_{d}(\bm{x})$.
We emphasize that the pre-defined $D$ domains do not need to match CTTA target domains. The main role of the WAD is to differentiate routers so the model can aggregate different visual features during continual TTA phases. Our BECoTTA is able to update relevant MODE modules with respect to the inputs and consistently achieves competitive performance even when the WAD and target domains are disjoint. (Please refer to Tab. 6 for details.)
Maximizing Domain-Expert Synergy. In cases where some domains share similar visual contexts (e.g., snow and fog), collaboration between domain experts can be beneficial. On the other hand, for unique scenes like night, it is advantageous to isolate domain features from others. That is, ensuring strong interdependence among various domains and experts is essential. To this end, we propose Domain-Expert Synergy loss based on the output from domain-adaptive routers. Let us consider $G_{d}^{i}(\bm{x})$ is assignment weight with specific domain $d$ from the $i$-th expert $A_{i}$, then $P(A_{i}|d)$ is obtained from all of the experts and domains in each MoDE layer. Then, we calculate $P(A_{i},d)$ using the Bayes’ theorem:

### v3

@@ Abstract
Continual Test Time Adaptation (CTTA) is required to adapt efficiently to continuous unseen domains while retaining previously learned knowledge. However, despite the progress of CTTA, it is still challenging to deploy the model with improved forgetting-adaptation trade-offs and efficiency. In addition, current CTTA scenarios assume only the disjoint situation, even though real-world domains are seamlessly changed. To address these challenges, this paper proposes BECoTTA, an input-dependent and efficient modular framework for CTTA. We propose Mixture-of-Domain Low-rank Experts (MoDE) that contains two core components: (i) Domain-Adaptive Routing, which helps to selectively capture the domain-adaptive knowledge with multiple domain routers, and (ii) Domain-Expert Synergy Loss to maximize the dependency between each domain and expert. We validate that our method outperforms multiple CTTA scenarios, including disjoint and gradual domain shits, while only requiring $\sim$98fewer trainable parameters. We also provide analyses of our method, including the construction of experts, the effect of domain-adaptive experts, and visualizations. Project Page: https://becotta-ctta.github.io/.
@@ 1 Introduction
Figure 1: BECoTTA and BECoTTA+ achieve superior 10-round average IoU and parameter/memory efficiency against strong CTTA baselines on the CDS-hard scenario.
@@
Figure 2: Comparison of TTA process with other SoTA models. We compare the existing models (Wang et al., 2020, 2022a; Song et al., 2023) and denote activated modules as yellow during CTTA process. In particular, CoTTA adopts the mean-teacher architecture and updates the entire model. TENT (Wang et al., 2020) and EcoTTA (Song et al., 2023) update only a few parameter-efficient modules in the model. However, they achieve suboptimal performance with forgetting. Meanwhile, our BECoTTA updates only MoDE layers for efficient and rapid adaptation while preserving previous knowledge.
@@
Beyond the isolated transferability of traditional TTA approaches on the stationary target domain, Continual Test-Time Adaptation (CTTA) (Niu et al., 2023; Song et al., 2023; Lee et al., 2023a) has been increasingly investigated in recent years, whose goal is to continuously adapt to multiple unseen domains arriving in sequence. Solving the problem of CTTA is crucial because it is closely related to real-world scenarios. For example, let us assume that a vision model in an autonomous vehicle is designed to understand road conditions and objects, including pedestrians, vehicles, traffic signs, etc. The agent will encounter different environments over time, depending on changes in weather, time of day, and location. Then, the model should rapidly and continuously adapt to these unseen environments while retaining the domain knowledge learned during adaptation as it may re-encounter prior domains in the future.
@@
Therefore, continual TTA approaches need to address the following key challenges: (i) forgetting-adaptation trade-off: retaining previous domain knowledge while learning new domains often limits the model’s plasticity, hindering its ability to learn and adapt to new data, and (ii) computational efficiency: since CTTA models are often assumed to be embedded in edge devices, efficient adaptation is significant. However, existing CTTA methods overlook computational efficiency by updating heavy teacher and student models (Wang et al., 2022a) or achieved suboptimal convergence due to updating only a few parts of modules (Wang et al., 2020; Niu et al., 2023; Gan et al., 2023; Gao et al., 2022).
@@
To tackle these critical issues, this paper proposes Input-dependent Online Blending of Experts for Continual Test-Time Adaptation (BECoTTA) by introducing a surprisingly efficient yet effective module, named Mixture-of-Domain Low-rank Experts (MoDE), atop each backbone block. Our BECoTTA method consists of two key components: (i) Domain Adaptive Routing and (ii) Domain-Expert Synergy Loss. We first propose Domain Adaptive Routing that aims to cluster lightweight low-rank experts (i.e., MoDE modules) with relevant domain knowledge. Next, based on the assignment of domain adaptive routers, we propose Domain-Expert Synergy Loss to maximize mutual information between each domain and its corresponding expert. In the end, we facilitate cooperation and specialization among domain experts by ensuring strong dependencies. Our modular design allows for selective updates of multiple domain experts, ensuring the transfer of knowledge for each specific domain while preserving previously acquired knowledge. This approach also significantly improves memory and parameter efficiency through sparse updates.
@@
Furthermore, existing CTTA scenarios assume a disjoint change of test domains, where the model encounters a static domain per time step, but do not consider a gradual shift of domains, which is more common in the real world (e.g., seamless weather change like cloudy $\rightarrow$ rainy or afternoon $\rightarrow$ night). To further consider this realistic scenario, we additionally propose Continual Gradual Shifts (CGS) benchmark for CTTA, where the domain gradually shifts over time based on the domain-dependent sampling distribution, as illustrated in Fig. 3 top left. This scenario is more advanced than an existing CTTA problem as it demands the model to appropriately adapt each of the input instances, without relying on any implicit guidance from the dominant domain over a given time interval.
@@
We compare our proposed method with strong baselines, including SAR (Niu et al., 2023), DePT (Gao et al., 2022), VDP (Gan et al., 2023), and EcoTTA (Song et al., 2023), on multiple CTTA scenarios and our suggested CGS benchmark. Our BECoTTA achieves +2.1%p and +1.7%p IoU enhancement respectively on CDS-Hard and CDS-Easy scenarios, by utilizing 95% and 98% fewer parameters used by CoTTA (Wang et al., 2022a). Furthermore, we propose BECoTTA+, which is initialized by the source augmentation dataset. BECoTTA+ shows increased performance by +16.8%p compared to EcoTTA (initialized fairly), utilizing a similar number of parameters on the CDS-Hard scenario.
@@ 2 Related Works
Continual Test-Time Adaptation (CTTA) (Wang et al., 2022a; Gan et al., 2023; Niu et al., 2023; Song et al., 2023) assumes that target domains are not fixed but change continuously in an online manner. TENT (Wang et al., 2020) is one of the pioneering works, which activates only BatchNorm layers to update trainable affine transform parameters. CoTTA (Wang et al., 2022a) introduces a teacher-student framework, generating pseudo-labels from the teacher model, and updating it using consistency loss. EcoTTA (Song et al., 2023) utilizes meta-networks and self-distilled regularization while considering memory efficiency. DePT (Gao et al., 2022) integrates visual prompts to efficiently adapt target domains and bootstraps the source representation. However, existing methods often suffer from subordinate convergence, as they rely on a shared architecture to adapt the test data without considering the correlation between different domains. On the other hand, our BECoTTA introduces a modularized MoE-based architecture where each expert captures domain-adaptive knowledge, and the model transfers only a few related experts for the adaptation of a new domain.
Moreover, recent works (Song et al., 2023; Niu et al., 2022; Lim et al., 2023; Choi et al., 2022a; Liu et al., 2021a; Adachi et al., 2022; Jung et al., 2023; Lee et al., 2023b) allow for a slight warm-up using the source dataset before deploying the model to the CTTA scenario. In particular, TTA-COPE (Lee et al., 2023b) performs pretraining with labeled source datasets in a supervised manner. EcoTTA (Song et al., 2023) also allows warm-up to initialize their meta-network. Note that these methods still assume source-free training during test-time adaptation, which means that this setup adheres to the assumptions of CTTA.
@@
Mixture-of-Experts (MoE) (Shazeer et al., 2017; Fedus et al., 2022; Zuo et al., 2021; Wang et al., 2022b) introduces $N$ parallel experts consisting of the feedforward network with router modules and sparsely activates a few experts based on their sampling policies. Adamix (Wang et al., 2022b) introduces efficient fine-tuning with Mixture-of-Adapters to learn multiple views from different experts. THOR (Zuo et al., 2021) proposes a new stochastic routing function to prevent inefficiency with routers. Similarly, Meta DMoE (Zhong et al., 2022) adopts the MoE architecture as a teacher model and distills their knowledge to unlabeled target domains, but does not consider continuous adaptation. In short, to the best of our knowledge, the feasibility of MoE structures is underestimated in the field of CTTA.
@@
Recently, a few continual learning appraoches (Koh et al., 2021; Bang et al., 2021, 2022; Aljundi et al., 2019) have discussed Blurry Continual Learning (Blurry-CL) to better reflect real-world scenarios, beyond the standard CL setting. Blurry-CL assumes that for each sequential task, a majority class exists, and other classes outside the majority class may also overlap and appear. The most renowned scenario setup is a Blurry-M (Aljundi et al., 2019), where the majority class occupies 100-M%, and the remaining classes are randomly composed of M%. Although this benchmark handles an overlapping situation, it may not cover practical situations where the domain evolves gradually in CTTA. Therefore, we propose a new benchmark that simulates real-world continual TTA scenarios with a gradual change of domains over time.
@@ 3 Input-dependent Online Blending of Experts for Continual Test-time Adaptation
We first define the problem statement for Continual Test-Time Adaptation (CTTA) in Sec. 3.1. Next, we introduce our proposed CTTA method, BeCoTTA, containing Mixture of Domain low-rank Experts (MoDE) and domain-expert synergy loss in Secs. 3.2 and 3.3. Then, we describe the overall optimization process during CTTA in Sec. 3.4.
@@ 3.1 Problem Statement
Continual Test-time Adaptation (CTTA) aims to adapt the pre-trained source model $\bm{f_{s}}$ to continuously changing target domains, formulated as a task sequence $\bm{X_{t}}=\{X_{t}^{1},X_{t}^{2},..X_{t}^{c},\cdots\}$. The main assumption of CTTA includes that (i) we should not access the source dataset after deploying the model to the test time, and (ii) adaptation needs to be done online and in an unsupervised manner. For semantic segmentation tasks, CTTA aims to predict the softmax output $\hat{y_{c}}=f_{c}(\bm{x}_{t}^{c})$ in the target domain $c$. $\bm{x}_{t}^{c}$ is sampled from $X_{t}^{c}$, which will be represented by $\bm{x}$ in the following sections for brevity.
Figure 3: The overview of BECoTTA. We propose a novel CTTA framework for dynamic real-world scenarios, including disjoint and gradual shifts of domains. When the model receives a target domain input $\bm{x_{t}}$ at timestep $t$, the Domain Discriminator (DD) first estimates a pseudo-domain label $d$. Based on estimated pseudo-labels, the domain router $G_{d}$ processes the input to specific experts containing domain-specific information by minimizing Domain-Expert Synergy Loss $\Theta(D;A)$. Finally, we obtain a domain-adaptive representation $h_{d}(\bm{x})$, addressing downstream tasks in test-time.
3.2 Domain-Augmented Initialization.
Source Domain Augmentation (SDA). Most CTTA methods use a pre-trained frozen backbone, which contains domain bias from the source domain. Due to the predominance of the source domain, this bias impedes the effective transfer of domain-adaptive knowledge in continuous scenarios. To mitigate this limitation, we define $D$ proxy domains (e.g., brightness, darkness, blur, etc.) and augment the source dataset to proxy domains, similar to EcoTTA (Song et al., 2023). For this Source Domain Augmentation (SDA), we utilize pre-trained style-transfer (Jiang et al., 2020) or simple transformations (Buslaev et al., 2020). Through SDA, we acquire domain-specific knowledge before deploying TTA. This process is done only once when constructing the source dataset.
Robustness to SDA. We emphasize that these pre-defined domains do not need to match CTTA target domains. The primary role of SDA is to differentiate routers so that the model can aggregate different visual features during the continual TTA phases. Our BECoTTA+ is able to update relevant MoDE modules with respect to the inputs and consistently achieves competitive performance even when the SDA and target domains are disjoint. (Please refer to Tab. 20 and Appendix (Sec. D) for more details.) For a fair comparison with other CTTA methods, we also verify that random and source-domain-only initializations work well with BECoTTA. More details are provided in Appendix (Tab. 17).
@@ 3.3 Mixture-of-Domain Low-rank Experts (MoDE)
The design of Low-rank Experts. For the efficient process during CTTA, we adopt the Sparse Mixture-of-Experts (SMoE) module with a top-k routing policy (Shazeer et al., 2017). Each MoE layer consists of the router $G$ and a set of $N$ lightweight experts, $A_{1}$, $A_{2}$, …, $A_{N}$, where each $A_{i}$ is parameterized by $\bm{W}_{i}^{down}\in\mathbb{R}^{dim\times r}$ and $\bm{W}_{i}^{up}\in\mathbb{R}^{r\times dim}$. Here, $r$ denotes the rank, and $dim$ denotes the embedding dimension of each ViT block. If $A_{i}$ is activated, it maps the input $\bm{x}$ into the low-dimensional space through the projection operation with $\bm{W}_{i}^{down}$. Next, after regularizing the features with a non-linear activation function $\sigma(\cdot)$, it recovers the features to the original dimension using $\bm{W}_{i}^{up}$:
@@
Based on the $G_{d}(\bm{x})$, we selectively update the activated experts associated with the specific domain, inherently isolating them from irrelevant domain knowledge while adapting new ones. In the end, the output of the MoDE layer $h_{d}(x)$ aggregates the domain-adaptive features as follows:
@@
This trainable clustering approach allows the model to activate its own set of experts, who are specialized in specific domain knowledge. Moreover, our multi-router-based design accelerates adaptation to the current domain, avoiding interference from knowledge transfer of unrelated domain features. Finally, we perform the skip connection operation with the original input $\bm{x}$: $\bm{x}\leftarrow\bm{x}+h_{d}(\bm{x})$.
Maximizing Domain-Expert Synergy. In cases where some domains share similar visual contexts (e.g., snow and fog), collaboration between domain experts can be beneficial. On the other hand, for unique scenes like night, it is advantageous to isolate domain features from others. That is, ensuring strong interdependence among various domains and experts is essential. To this end, we propose Domain-Expert Synergy loss based on the output from domain-adaptive routers. Let us consider $G_{d}^{i}(\bm{x})$ as the assignment weight with specific domain $d$ of the $i$-th expert $A_{i}$, then $P(A_{i}|d)$ is obtained from all the experts and domains in each MoDE layer. Then, we calculate $P(A_{i},d)$ using Bayes’ theorem: